OpenStudy (anonymous):

Name the set(s) of numbers to which 1.68 belongs.

OpenStudy (ybarrap):

Rational 1.68 because 1.68 = 168/100 Real Number because \(-\infty \lt x\lt \infty\)

OpenStudy (ybarrap):

And \(x=1.68\) (above). Also, it is algebraic because 25 x-42=0 \(\implies x=1.68\)